Lora (of mass 43.3 kg) is an expert skier. She
starts at 2.9 m/s at the top of the lynx run,
which is 106 m above the bottom.
What is her final kinetic energy at the bottom of the ski run?
Answer in units of j (due in 3 hours)
1
Expert's answer
2012-03-29T13:12:19-0400
Lora (of mass m = 43.3 kg) is an expert skier. She starts at V = 2.9 m/s at the top of the lynx run, which is H = 106 m above the bottom.
What is her final kinetic energy at the bottom of the ski run? Answer in units of J.

The final kinetic energy can be represented as

E = E0 + E1,

where

E0 = mV²/2 = 43.3*2.9²/2 = 182.0765 J

is the initial kinetic energy and

E1 = mgH = 43.3*9.8*106 = 44980.04 J

is the potential energy that turns to kinetic while moving down. So,

E = 182.0765 + 44980.04 = 45162.1165 J.